a. To the nearest power of 10, what is the population of the United States? b. To the nearest power of 10, how old is the Earth is years? c. To the nearest power of 10, what is the surface area of the Earth?
Answer #1

Population of united states = 3.28*108

Age of Earth = 4.5*109 years

Surface area of Earth = 5.1*1014 m​​​​​​m2